OpenStudy (anonymous):

What is the value of -81^(1/4)

OpenStudy (anonymous):

this is really: \[\sqrt[4]{-81}\]

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

langrangeson, when i put it in my calculator it won't do it J3553, I got that when I put -81^.25 but -3times-3=9 9times-3=-27 -27times-3=81 not -81

OpenStudy (anonymous):

That's why I'm so confused

OpenStudy (anonymous):

-81^1/4 = (-1)^1/4 x (81)^1/4 now, -1 = i^2 so -81^1/4 = i^1/2 x 3 i^1/2 = 1/sqrt(2) x(1+i)

OpenStudy (anonymous):

What just happened??????

OpenStudy (anonymous):

Someone please help!

OpenStudy (anonymous):

calm down everything will be okay, do you still need help?

OpenStudy (anonymous):

Yes! Look at the third post.

OpenStudy (anonymous):

okay, this is what we will do, first we factor out the negative in the form of -1

OpenStudy (anonymous):

Ok, I just need to get it straight whether it's neg or pos 3, so lets do this

OpenStudy (anonymous):

OH! I think I get where you're going with this!

OpenStudy (anonymous):

right, so what i meant to type is :\[-1*\sqrt[4]{81}\]

OpenStudy (anonymous):

Ok! Thank you so much!

OpenStudy (anonymous):

now just do your thing from there